OpenStudy (anonymous):

a line perpendicular to the graph of 4x+y=7

OpenStudy (mathstudent55):

Write 4x + y = 7 in y = mx + b form: y = -4x + 7 The slope of the given line is -4, so the slope of a perpendicular is 1/4.

OpenStudy (unklerhaukus):

because the slopes of perpendicular lines multiply to give -1 \[m_1m_2=-1\]

OpenStudy (raden):

alternatiave : the line perpendicular with the line ax+by=c and passes through point (x1,y1) is bx - ay = b(x1) - a(y1)

OpenStudy (anonymous):

how would it be 1/4 wouldnt it just be -4?

OpenStudy (unklerhaukus):

\[m_1m_2=-1\] \[m_2=\frac{-1}{m_1}\] \[\qquad\qquad m_1=-4\] \[m_2=\frac{-1}{-4}=\frac14\]

OpenStudy (mathstudent55):

-4 is the slope of the given line. The slopes of perpendicular lines are negative reciprocals. That means that when you multiply the slopes together you get -1. Since ine slope is -4, the other one is 1/4 becasue -4 * (1/4) = -1

OpenStudy (mathstudent55):

Your question is simply a line perpendicular to a given line. There is an infinite number of lines perpendicular to a given line. If you stated that you're looking for a line perpendicular to 4x + y = 7 that passes through a specific point, then we can determine the equation of the specific line you are looking for. But sinc e you only mention a line perpendicular to a given line, all that can be done is to tell you what the slope of any perpendicular has to be.

OpenStudy (anonymous):

becuz we are looking for a line that is perpendicular to the 4x+y=7.. okay kind of confusing in the begining but after you explained i understood it. thank you

OpenStudy (anonymous):

for your time and i really appreciate it!!!!

OpenStudy (mathstudent55):

You are very welcome.